OpenStudy (ajanijones):

Janise rolled a number cube 50 times and recorded the results in this table. Outcome Number 1 7 2 9 3 8 4 8 5 6 6 12 What is the theoretical probability of the cube landing on a 6? What is the experimental probability of the cube landing on a 6? A. Theoretical - 12% Experimental - 24% B. Theoretical - 16.7% Experimental - 16.7% C. Theoretical - 16.7% Experimental - 24% D. Theoretical - 20% Experimental - 12%

OpenStudy (tkhunny):

Unfortunately, we don't ACTUALLY know enough about the "Number Cube". If it had said a "Standard Die", rather than the oddly-more-socially-acceptable "Number Cube", we would have had more information. Let's assume the Number Cube is a Standard Die, making all of 1-6 equally probable. p(6) = 1/6 Experimental Result = 12/50 Create decimal equivalents and you are done.

OpenStudy (ajanijones):

Ok that you I though something was wrong with not know enough. well thank you.

OpenStudy (tkhunny):

If you plan to IM me, you'll have to "FAN" me if you want a response in that environment. However, it's fine that I can't IM you, since I woudl MUCH rather communicate here, publicly and where LaTeX works. 12/50 = 0.24 = 24%. What's the question? You need to be fluent in this sort of transformation.

OpenStudy (ajanijones):

So my answer's A.

OpenStudy (tkhunny):

That is not correct. Take another shot at p(6) = 1/6.